# Not All Steps are Informative: On the Linearity of LLMs’ RLVR Training

### v1

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has become a central component of large language model (LLM) post-training. Unlike supervised fine-tuning (SFT), RLVR lets an LLM generate multiple candidate solutions and reinforces those that lead to a verifiably correct final answer. However, in practice, RLVR often requires thousands of training steps to reach strong performance, incurring substantial computation largely attributed to prolonged exploration. In this work, we make a surprising observation: during RLVR, LLMs evolve in a strongly linear manner. Specifically, both model weights and model output log-probabilities exhibit strong linear correlations with RL training steps. This suggests that RLVR predominantly amplifies trends that emerge early in training, rather than continuously discovering new behaviors throughout the entire optimization trajectory. Motivated by this linearity, we investigate whether future model states can be predicted from intermediate checkpoints via extrapolation, avoiding continued expensive training. We show that Weight Extrapolation produces models with performance comparable to standard RL training while requiring significantly less computation. Moreover, Logits Extrapolation consistently outperforms continued RL training on all four benchmarks by extrapolating beyond the step range where RL training remains stable.
@@
Tianle Wang1,2,3*  Zhongyuan Wu3,4*  Shenghao Jin3,4  Hao Xu3  Wei Chen3  Ning Miao1,2,3
@@ 1 Introduction
Figure 1: Linearity analysis for model weights and outputs during RLVR training. (a) and (b) show the distributions of $R^{2}$ for weight and token log-probabilities, respectively. Both distributions are concentrated around 0.9, indicating strong linearity. (c) plots the trajectories of four randomly selected weights, and (d) shows token log-probability changes at four example positions. The log-probabilities of “wait” and “but” increase over RL steps, suggesting more reflection and revision, whereas those of “earlier” and “alternatively” decrease, indicating reduced need for backtracking and branching.
@@
For weight linearity, we performed linear regression of weights at all checkpoints against training steps. We observe clear linear weight dynamics: as training progresses, the majority of weights change in a highly linear fashion. For example, as shown in Figure 1(c), weight #1, #6, and #9 exhibit near-linear increases, whereas parameter #0 decreases nearly linearly; Figure 1(c) shows the $R^{2}$ distribution of all weights, excluding those that are unchanged during training. We can see that more than 80% of the weights achieve $R^{2}>0.7$, indicating strong weight linearity.
We also performed controlled checkpoint-probing to analyze the linearity of model outputs. Specifically, we selected a set of queries and their solution trajectories as probes. For each checkpoint, we computed the log-probabilities of each token, conditional on all previous tokens. As shown in Figure 1(a) and 1(d), there is also a strong linear correlation between model predicted log-probabilities and training steps. We also observed linearity of other model outputs, such as logits, and intermediate activations, whose results are left in Appendix A.2.
@@ 3 Linearity of RLVR Training
Figure 2: Linearity consistency across diverse experimental setups. The $R^{2}$ scores consistently exceed 0.7 (dashed line) across various base models (e.g., DS-Qwen, DS-Llama), scale sizes (1.5B to 8B), and training algorithms (GSPO, Reinforce++, and GRPO). The high $R^{2}$ values for both token log-probabilities and weights indicate a robust linear relationship that persists across architectural and algorithmic configurations.
@@ 4 Accelerating RLVR with linearities
The linearities of RLVR training indicate that the weights and output at a certain step can be largely predicted by its training trajectory at earlier steps. As a result, we can speed up RLVR training by replacing standard training of some steps with linear extrapolation. In the following, we will describe direct extrapolation from two perspectives: Weight Extrapolation and Logits Extrapolation. We then introduce RL-Extra, an interactive training scheme that interleaves extrapolation with RL updates.
@@ 4.2.1 Logits Extrapolation
Figure 4: Accuracy comparison on AIME and LCB benchmarks. Logit Extrapolation yields consistent improvements over standard RL across all evaluated settings.

### v2

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has become a central component of large language model (LLM) post-training. Unlike supervised fine-tuning (SFT), RLVR lets an LLM generate multiple candidate solutions and reinforces those that lead to a verifiably correct final answer. However, in practice, RLVR often requires thousands of training steps to reach strong performance, incurring substantial computation largely attributed to prolonged exploration. In this work, we make a surprising observation: during RLVR, LLMs evolve in a strongly linear manner. Specifically, both model weights and model output log-probabilities exhibit strong linear correlations with RL training steps. This suggests that RLVR predominantly amplifies trends that emerge early in training, rather than continuously discovering new behaviors throughout the entire optimization trajectory. Motivated by this linearity, we investigate whether future model states can be predicted from intermediate checkpoints via extrapolation, avoiding continued expensive training. We show that Weight Extrapolation produces models with performance comparable to standard RL training while requiring significantly less computation. Moreover, Logits Extrapolation consistently outperforms continued RL training on mathematics and code benchmarks by extrapolating beyond the step range where RL training remains stable. Our code is available at https://github.com/Miaow-Lab/RLVR-Linearity
@@
Tianle Wang1,2,3*  Zhongyuan Wu3,4*  Shenghao Jin3,4  Hao Xu3  Wei Chen3  Ning Miao1,2
@@ 1 Introduction
Figure 1: Linearity analysis for model weights and outputs during RLVR training. (a) and (b) show the distributions of $R^{2}$ for weight and token log-probabilities, respectively. Both distributions are concentrated around 0.9, indicating strong linearity. (c) plots the trajectories of four randomly selected weights, and (d) shows token log-probability changes at four example positions.
@@
For weight linearity, we performed linear regression of weights at all checkpoints against training steps. We observe clear linear weight dynamics: as training progresses, the majority of weights change in a highly linear fashion. For example, as shown in Figure 1(c), weight #1, #6, and #9 exhibit near-linear increases, whereas parameter #0 decreases nearly linearly; Figure 1(a) shows the $R^{2}$ distribution of all weights, excluding those that are unchanged during training. We can see that more than 80% of the weights achieve $R^{2}>0.7$, indicating strong weight linearity.
We also performed controlled checkpoint-probing to analyze the linearity of model outputs. Specifically, we selected a set of queries and their solution trajectories as probes. For each checkpoint, we computed the log-probabilities of each token, conditional on all previous tokens. As shown in Figure 1(b) and 1(d), there is also a strong linear correlation between model predicted log-probabilities and training steps. The log-probabilities of “wait” and “but” increase over RL steps, suggesting more reflection and revision, whereas those of “earlier” and “alternatively” decrease, indicating reduced need for backtracking and branching. We also observed linearity of other model outputs, such as logits, and intermediate activations, whose results are left in Appendix A.2.
@@ 3 Linearity of RLVR Training
Figure 2: Linearity consistency across diverse experimental setups. The $R^{2}$ scores consistently exceed 0.7 (dashed line) across various base models (e.g., DS-Qwen, DS-Llama), scale sizes (1.5B to 8B), and training algorithms (GSPO, Reinforce++, and GRPO).
@@ 4 Accelerating RLVR with linearities
Figure 4: Overall framework of three proposed methods to accelerate RLVR.
The linearities of RLVR training indicate that the weights and output at a certain step can be largely predicted by its training trajectory at earlier steps. As a result, we can speed up RLVR training by replacing standard training of some steps with linear extrapolation. In the following, we will describe direct extrapolation from two perspectives: Weight Extrapolation and Logits Extrapolation. We then introduce RL-Extra, an interactive training scheme that interleaves extrapolation with RL updates. The extrapolation framework is illustrated in Figure 4.
@@ 4.2.1 Logits Extrapolation
Figure 5: Accuracy comparison on AIME and LiveCodeBench (LCB) benchmarks. Logit Extrapolation yields consistent improvements over standard RL across all evaluated settings.